find the slope and the y-intercept of the line. Write the fractional answer in lowest terms. y = -6/5x - 5/6
the slope is ?
the y-intercept is (0,?)

Answer by mananth(16949)

y = -6/5x - 5/6
compare with y= mx+b
m= slope b= y intercept
m=-6/5
b=-5/6